Sonja Cc is a small firm that make outfits for customer tailored for their functions, in My 2021 she was ordered for a baby shower (BS)

The details for Job BS was as follows:

Direct materials: 12 units at N$124

Direct labour: 20 hours for department A at N$25 per hour

10 hours for department B at N$30 per hour

12 hours for department C at N$27.50 per hour


The firm allocates manufacturing overheads budgeted at N$2500 in department A, N$2000 in department B and N$3000 in department C.

The firm expects a profit of 55% on total cost BS


Required:

1.Calculate the cost of a baby shower BS

2.Calculate the selling price of a baby shower BS

3.If the number of outfits are the same for each type of function, which function should Sonja choose to produce and why.


1
Expert's answer
2022-03-29T12:09:22-0400

Actual cost (without profit) ="12\times124+20\times25+10\times30+12\times27.5=2618"


selling price of baby shower

"=\frac{100+55}{100}\times2618\\=\frac{155}{100}\times2618\\=4,057.9"


Sonja should choose department B because it allocates the least overhead cost.